please help me solve this problem.
two missiles are 450 miles apart and are moving toward each other.one missile travels 100 mph faster than the other.in 30 minute4s the missiles collide.what was the rate of speed of each missile?

two ships are 900 miles apart and moving toward each other.one ship travels twice as fast as the other.in 6 hours the ship pass each other what is the rate of speed of each ship?
armies A and B are situated 26,400 feet apart.Army A fires a missile traveling at 8600 feet per second at Army B.Two seconds later,Army B shoots an antibalistic missile at Army A's missile and hits it 400 feet from Army B's position.how fast is Army B's missile moving?

Answer by stanbon(75887)

two missiles are 450 miles apart and are moving toward each other.one missile travels 100 mph faster than the other.in 30 minutes the missiles collide.what was the rate of speed of each missile?
1st missile DATA:
time = (1/2)hr ; rate = x+100 mph ; distance = t*r = (1/2)(x+100) miles
------------
2nd missile DATA:
time = (1/2)hr ; rate = x mph ; distance = t*r = (1/2)x miles
------------
Equation:
distance + distance = 450 miles
(1/2)(x+100) + (1/2)x = 450
x + 50 = 450
x = 400
--
Rate of the 1st missile: x+100 = 500 mph
Rate of the 2nd millile: x = 400 mph
====================

two ships are 900 miles apart and moving toward each other.one ship travels twice as fast as the other.in 6 hours the ship pass each other what is the rate of speed of each ship?
---
Same procedure as the missile problem.
---

armies A and B are situated 26,400 feet apart. Army A fires a missile traveling at 8600 feet per second at Army B.Two seconds later,Army B shoots an antibalistic missile at Army A's missile and hits it 400 feet from Army B's position.how fast is Army B's missile moving?
A-missile DATA:
rate = 8600 ft/sec ; distance = 26000 ft; time = 26000/8600 = 3.0233 sec
-----
B-missile DATA:
distance = 400 ft; time = 1.0233 sec ; rate = 400/1.0233 = 390.91 ft/sec
--------------
Cheers,
Stan H.